# Uni3C: Unifying Precisely 3D-Enhanced Camera and Human Motion Controls for Video Generation

### v1

@@ Abstract
Camera and human motion controls have been extensively studied for video generation, but existing approaches typically address them separately, suffering from limited data with high-quality annotations for both aspects. To overcome this, we present Uni3C, a unified 3D-enhanced framework for precise control of both camera and human motion in video generation. Uni3C includes two key contributions. First, we propose a plug-and-play control module trained with a frozen video generative backbone, PCDController, which utilizes unprojected point clouds from monocular depth to achieve accurate camera control. By leveraging the strong 3D priors of point clouds and the powerful capacities of video foundational models, PCDController shows impressive generalization, performing well regardless of whether the inference backbone is frozen or fine-tuned. This flexibility enables different modules of Uni3C to be trained in specific domains, i.e., either camera control or human motion control, reducing the dependency on jointly annotated data. Second, we propose a jointly aligned 3D world guidance for the inference phase that seamlessly integrates both scenic point clouds and SMPL-X characters to unify the control signals for camera and human motion, respectively. Extensive experiments confirm that PCDController enjoys strong robustness in driving camera motion for fine-tuned backbones of video generation. Uni3C substantially outperforms competitors in both camera controllability and human motion quality. Additionally, we collect tailored validation sets featuring challenging camera movements and human actions to validate the effectiveness of our method.
[…]
1 Introduction
Figure 1: Given a single-view image across various domains (e.g., real-world, text-to-image, animation), we first extract the monocular depth and focal length of it via Depth-Pro [6] and then achieve relative point clouds. Then, the proposed Uni3C can generate impressive videos under arbitrary camera trajectories (a), human motion characters (SMPL-X [36]), or both of these conditions (b). (c) Uni3C further supports the camera-controlled motion transfer.
@@ 1 Introduction
Recent advancements in foundational video diffusion models (VDMs) [5, 64, 7, 27, 46, 26, 54] have unlocked unprecedented capabilities in creating dynamic and realistic video content. A significant challenge in this field is achieving controllable video generation, a feature with broad applications in virtual reality, film production, and interactive media. In this paper, we focus on two aspects of controllable video generation: camera control [17, 63, 69, 2, 66, 44] and human motion control [21, 57, 62, 72, 51, 22]—both of which are critical and interdependent in real-world scenarios.
@@
Recent pioneering works have extensively studied controlling camera trajectories for VDMs through explicit conditions like Plücker ray [17, 2, 69, 18, 30], point clouds [66, 44, 28, 12, 40]. Concurrently, controllable human animation also attracted a lot of attention based on poses [21, 62, 51, 22] or SMPL characters [72, 75, 73]. Despite these advancements, several challenges remain: 1) Most approaches deeply hack the inherent capacities of VDMs, which have been trained with domain-specific data and conditions, inevitably undermining the generalization to handle out-of-distribution scenarios. 2) Very few works investigate the joint control of both camera trajectories and human motions. This requires diverse camera trajectories in human-centric videos with high-quality annotations [57], which are often expensive to obtain. 3) There is a lack of explicit and synchronized guidance that incorporates strong 3D-informed priors to concurrently control both camera movements and human motions. Relying on separate conditions, like point clouds and SMPL, struggles to represent physically reasonable interactions and positional relations between characters and environments.
@@
To address these challenges, we present Uni3C, a novel framework that Unifies precise 3D-enhanced camera and human motion Controls for video generation as shown in Figure 1 via two key innovations. Firstly, we propose the insight that controlling camera trajectories for powerful foundational VDMs can be achieved by lightweight, trainable modules with rich informed guidance and reasonable training strategies. By avoiding hacking the underlying capacities of VDMs, our camera-controlling model can be directly generalized to versatile downstream tasks rather than costly joint training and extensive Structure from Motion (SfM) labeling [47, 68, 29]. Secondly, we claimed that camera and human motion controls are inherently interdependent. Therefore, we propose to align their conditions into a global 3D world during the inference phase, enabling 3D consistent generation across both domains.
@@
Formally, Uni3C is built upon the foundational VDM—Wan2.1 [54]. For the control of camera trajectories, we propose PCDController, a plug-and-play control module with only 0.95B parameters (compared to 12B of Wan2.1) that operates on unprojected 3D point clouds derived from monocular depth estimation [6]. Thanks to the rich geometric priors of point clouds, PCDController is capable of fine-grained camera control, even when trained on constrained multi-view images and videos with a frozen backbone. Furthermore, PCDController can be compatible with fine-tuned VDM backbones for versatile downstream tasks. This surprising factor supports domain-specific training, i.e., camera and human motion modules can be trained independently without jointly annotated data. For the global 3D world guidance, we align scenic point clouds (for camera control) and SMPL-X characters [36] (for human animation) into the same world-coordinate space via the rigid transformation [52], while the 2D keypoints [61] bridge the relation of two presentations. Note that our alignment enables complicated motion transfer, covering disparate characters, positions, and viewpoints as verified in the last row of Figure 1.
@@
Extensive experiments validate the efficacy of Uni3C. To evaluate the remarkable generalization of PCDController, we collect an out-of-distribution test set across different domains, where each image has four different camera trajectories. For the joint controllability, we build a comprehensive test set of in-the-wild human videos. GVHMR [48] is used to extract SMPL-X as the condition, while three complex and random camera trajectories are assigned for each video. VBench++ [23] is employed to verify the overall performance of our method. Uni3C significantly outperforms other competitors, both quantitatively and qualitatively.
@@
2 Related Work
Camera Control for VDMs.
@@ 2 Related Work
Controlling camera trajectories in video generation has garnered significant attention in recent times. Some works focused on injecting camera parameters into VDMs to achieve camera controllability [58, 17, 3, 2, 69, 18, 30], typically utilizing the Plücker ray presentation. For instance, VD3D [3] designed a tailored framework for Diffusion Transformer (DiT) [38], while AC3D [2] further emphasized the generalization with less trainable parameters. Moreover, DimensionX [50] further decoupled the spatial and temporal control with different LoRAs [20]. Despite the progress made by these methods, they cannot control the camera movements precisely, particularly when the case is beyond the training domains with an unknown metric scale. Thus, other recent works have employed point cloud conditions through both training-based [66, 44, 28, 12, 40] and training-free [19, 65] manners. However, these methods fail to accommodate the generalized model design and training strategy to handle imperfect point clouds or out-of-distribution data, especially in scenarios involving humans or animals.
Unified Control for VDMs.
@@
Recent works have unified multiple conditions to guide video generation [58, 12, 13, 57, 10, 14, 70]. MotionCtrl [58] integrated the camera and object motion controls through separate pose and trajectory injections. Subsequently, researchers further explored the presentation of conditions, such as point trajectories [12], point tracking [13, 14], and 3D-aware signals [10]. Humanvid [57] first unified the control of human scenes, while VidCraft3 [70] further considered the lighting control. Although these pioneering approaches achieved promising performance, they often rely heavily on joint training with various conditions and well-labeled datasets. Furthermore, there has been limited discussion on unified control within foundational VDMs that exceed 10B parameters. Our work offers a solution to address these issues: unifying existing models for different downstream tasks without the need for costly fine-tuning or fully annotated datasets. This strategy is particularly well-suited for large VDMs, enabling models to focus more on enhancing performance within their specific domains.
3 Preliminary: Video Diffusion Models
@@
We briefly review VDMs and Wan2.1 [54] in this section as preliminary knowledge. VDMs are mainly based on the latent diffusion model [45], modeling the conditional distribution $p(z_{0}|c_{txt},c_{img})$, where $z_{0}$ indicate clean video latent features encoded by 3D-VAE; $c_{txt},c_{img}$ denote the text condition for Text-to-Video (T2V) and the optional image condition for Image-to-Video (I2V), respectively. The training of VDM involves reversing the diffusion process by the noise estimator $\epsilon_{\theta}$ as:
$\min_{\theta}\mathbb{E}_{z_{0},t,\epsilon,c_{txt},c_{img}}[\|\epsilon_{\theta}(z_{t},t,c_{txt},c_{img})-\epsilon\|^{2}],$ (1)
@@
where $\epsilon\sim\mathcal{N}(0,I)$ indicates Gaussian noise; timestep $t\in[0,T_{max}]$; $z_{t}$ is the intermediate noisy latent state of timestep $t$. Recently, most VDMs have employed Flow Matching (FM) [33] as the improved diffusion process with faster convergence and more stable training. Based on the ordinary differential equations (ODEs), FM formulates the linear interpolation between $z_{0}$ and $z_{1}$, i.e., $z_{t}=tz_{1}+(1-t)z_{0}$, where $t\in[0,1]$ is sampled from the logit-normal distribution. The velocity prediction $v_{\theta}$ can be written as:
$\min_{\theta}\mathbb{E}_{z_{0},t,\epsilon,c_{txt},c_{img}}[\|v_{\theta}(z_{t},t,c_{txt},c_{img})-v_{t}\|^{2}],$ (2)
@@
where the ground truth velocity denotes $v_{t}=\frac{dz_{t}}{dt}=z_{1}-z_{0}$. Additionally, recent foundational VDMs [7, 64, 26, 54] are built with DiT [38] to achieve more capacities for video generation.
@@
Wan2.1 [54] is an open-released, large-scale VDM with DiT architecture trained with flow matching [33]. umT5 [11] is utilized as the multi-language text encoder to inject textual features into Wan2.1 through cross-attention. For image-to-video, Wan-I2V further incorporates features from CLIP’s image encoder [42] to improve the results. Uni3C is primarily designed for Wan-I2V with 14B parameters, but we empirically find that it is compatible with the Wan-T2V version as verified in Section 5.4, showing convinced generalization of PCDController.
4 Method
@@
Figure 2: The overview pipeline of PCDController. PCDController is built as a lightweight DiT trained from scratch. We first obtain the point clouds via the monocular depth extracted from the first view. Then, the point clouds are warped and rendered into the video $V_{pcd}$. The input conditions for PCDController comprise rendered $V_{pcd}$, Plücker ray $\mathbf{P}$, and the noisy latent $z_{t}$. Note that only the PCDController and camera encoder are trainable in our framework.
@@
Given a reference view $I_{img}\in\mathbb{R}^{3\times h\times w}$, camera trajectories $\{c_{cam}^{i}\}_{i=1}^{N}$ of $N$ target views, and textual condition, Uni3C uses PCDController to produce the target video $\{V_{tar}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 3\times h\times w}$ under specified camera trajectories. This can be formulated as $p(z_{0}|c_{txt},c_{img},c_{cam})$, where $z_{0}$ indicates the clean latent video feature, $c_{txt},c_{img}$ are textual features from umT5 and image latent condition encoded from $I_{img}$, respectively. We show the overview pipeline of PCDController in Figure 2, which is a core component of Uni3C for generalized camera control (Section 4.1) across diverse downstream tasks. In this work, we focus on human animation (Section 4.2). Subsequently, we introduce the global 3D world guidance illustrated in Figure 5 to unify both camera and human characters into a consistent 3D space for inference (Section 4.3).
4.1 PCDController with 3D Geometric Priors
@@
Following AC3D [2], PCDController is designed within a simplified DiT module rather than copying modules and weights from the main backbone as ControlNet [67]. To preserve the generalization of Wan-I2V, we follow the insight of training as few parameters as possible once the effective camera control has been achieved. Formally, we reduce the hidden size of PCDController from 5120 to 1024, while zero-initialized linear layers are used to project hidden size back to 5120 before being added to Wan-I2V. Moreover, as investigated in [2, 30], VDMs mainly determine camera information through shallow layers. Thus, we only inject camera-controlling features into the first 20 layers of Wan-I2V to further simplify the model. Additionally, we discard the textual condition for PCDController to alleviate intractable hallucination and remove all cross-attention modules. In this way, the overall number of trainable parameters for PCDController is reduced to 0.95B, a significant reduction compared to Wan-I2V (14B).
@@
In contrast to merely utilizing Plücker ray as the camera embedding [2, 30], we incorporate much stronger 3D geometric priors to compensate the model simplification, i.e., videos $\{V_{pcd}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 3\times h\times w}$ rendered from point clouds under given camera trajectories. Specifically, we first use Depth-Pro [6] to extract the monocular depth map from the reference view. We then align this depth map into a metric representation using SfM annotations [47] or multi-view stereo [8]. Following [9], we employ RANSAC to derive the rescale and shift coefficients, preventing the collapse of constant depth outcomes. Subsequently, the point clouds $X_{pcd}\in\mathbb{R}^{hw\times 3}$ are got by unprojecting all 2D pixels from $I_{img}$ into the world coordinate via its metric depth $\hat{D}_{img}$ as follows:
$X_{pcd}(x)\simeq R_{c\rightarrow w}\hat{D}_{img}(x)K^{-1}\hat{x},$ (3)
where $x$ denotes the 2D coordinates of $I_{img}$, while $\hat{x}$ refers to the homogeneous coordinates; $K,R_{c\rightarrow w}$ mean the intrinsic and extrinsic cameras of the reference view, respectively. After that, we render $\{V_{pcd}^{i}\}_{i=2}^{N}$ for the remaining ($N-1$) views by PyTorch3D according to their respective camera intrinsics and extrinsics. Note that the first rendering corresponds to the reference image, i.e., $V_{pcd}^{1}=I_{img}$, to confirm the identity. We apply $V_{pcd}$ to the 3D-VAE of Wan2.1 to achieve $c_{pcd}$ as the point latent condition. To further handle the significant viewpoint changes, which may extend beyond the point clouds’ visibility of the first frame, PCDController also includes Plücker ray embedding [60], $\{\mathbf{P}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 6\times h\times w}$, as the auxiliary condition. $\{\mathbf{P}^{i}\}_{i=1}^{N}$ is encoded by a small camera encoder, comprising casual convolutions and a 4-8-8 downsampling factor to keep the same sequential length as 3D-VAE outputs. Therefore, the distribution modeling of PCDController can be written as $p(z_{0}|c_{txt},c_{img},c_{pcd},\mathbf{P})$.
Figure 3: Results of PCDController with imperfect point clouds. Benefiting from the well-preserved capacity from VDM, PCDController enjoys robust generation with inferior point clouds.

### v2

@@ Abstract.
Camera and human motion controls have been extensively studied for video generation, but existing approaches typically address them separately, suffering from limited data with high-quality annotations for both aspects. To overcome this, we present Uni3C, a unified 3D-enhanced framework for precise control of both camera and human motion in video generation. Uni3C includes two key contributions. First, we propose a plug-and-play control module trained with a frozen video generative backbone, PCDController, which utilizes unprojected point clouds from monocular depth to achieve accurate camera control. By leveraging the strong 3D priors of point clouds and the powerful capacities of video foundational models, PCDController shows impressive generalization, performing well regardless of whether the inference backbone is frozen or fine-tuned. This flexibility enables different modules of Uni3C to be trained in specific domains, i.e., either camera control or human motion control, reducing the dependency on jointly annotated data. Second, we propose a jointly aligned 3D world guidance for the inference phase that seamlessly integrates both scenic point clouds and SMPL-X characters to unify the control signals for camera and human motion, respectively. Extensive experiments confirm that PCDController enjoys strong robustness in driving camera motion for fine-tuned backbones of video generation. Uni3C substantially outperforms competitors in both camera controllability and human motion quality. Additionally, we collect tailored validation sets featuring challenging camera movements and human actions to validate the effectiveness of our method. Codes are released at https://github.com/alibaba-damo-academy/Uni3C.
Camera control, Human animation, Video generation, Generative models
submissionid: 1200journal: TOGccs: Networks Network algorithmsccs: Computing methodologies Computer vision
1. Introduction
Recent advancements in foundational video diffusion models (VDMs)
@@ 1. Introduction
(Blattmann et al., 2023; Yang et al., 2025; Brooks et al., 2024; Kuaishou, 2024; RunwayML, 2024; Kong et al., 2024; Wang et al., 2025a) have unlocked unprecedented capabilities in creating dynamic and realistic video content. A significant challenge in this field is achieving controllable video generation, a feature with broad applications in virtual reality, film production, and interactive media. In this paper, we focus on two aspects of controllable video generation: camera control (He et al., 2025a; Yang et al., 2024; Zheng et al., 2025a; Bahmani et al., 2025a; Yu et al., 2024; Ren et al., 2025) and human motion control (Hu, 2024; Wang et al., 2024a; Xu et al., 2024b; Zhou et al., 2024; Tan et al., 2025; Hu et al., 2025)—both of which are critical and interdependent in real-world scenarios.
@@
Recent pioneering works have extensively studied controlling camera trajectories for VDMs through explicit conditions like Plücker ray (He et al., 2025a; Bahmani et al., 2025a; Zheng et al., 2025a; He et al., 2025b; Liang et al., 2024), point clouds (Yu et al., 2024; Ren et al., 2025; Li et al., 2025b; Feng et al., 2025; Popov et al., 2025). Concurrently, controllable human animation also attracted a lot of attention based on poses (Hu, 2024; Xu et al., 2024b; Tan et al., 2025; Hu et al., 2025) or SMPL characters (Zhou et al., 2024; Zhu et al., 2024; Zhou et al., 2025b). Despite these advancements, several challenges remain: 1) Most approaches deeply hack the inherent capacities of VDMs, which have been trained with domain-specific data and conditions, inevitably undermining the generalization to handle out-of-distribution scenarios. 2) Very few works investigate the joint control of both camera trajectories and human motions. This requires diverse camera trajectories in human-centric videos with high-quality annotations (Wang et al., 2024a), which are often expensive to obtain. 3) There is a lack of explicit and synchronized guidance that incorporates strong 3D-informed priors to control both camera movements and human motions concurrently. Relying on separate conditions, like point clouds and SMPL, struggles to represent physically reasonable interactions and positional relations between characters and environments, resulting in conflicting guidance and suboptimal outcomes.
@@
To address these challenges, we present Uni3C, a novel framework that Unifies precise 3D-enhanced camera and human motion Controls for video generation as shown in Figure 1 via two key innovations. Firstly, we propose the insight that controlling camera trajectories for powerful foundational VDMs can be achieved by lightweight, trainable modules with rich informed guidance and reasonable training strategies. By avoiding hacking the underlying capacities of VDMs, our camera-controlling model can be directly generalized to versatile downstream tasks rather than costly joint training and extensive Structure from Motion (SfM) labeling (Schönberger et al., 2016; Zhao et al., 2022; Li et al., 2025a). Secondly, we claimed that camera and human motion controls are inherently interdependent. Therefore, we propose to align their conditions into a global 3D world during the inference phase, enabling 3D consistent generation across both domains.
Formally, our Uni3C is built upon the foundational VDM—Wan2.1
@@
 (Wang et al., 2025a). For the control of camera trajectories, we propose PCDController, a plug-and-play control module with only 0.95B parameters (compared to 14B of Wan2.1) that operates on unprojected 3D point clouds derived from monocular depth estimation (Bochkovskii et al., 2024). Thanks to the rich geometric priors of point clouds, PCDController is capable of fine-grained camera control, even when trained on constrained multi-view images and videos with a frozen backbone. Furthermore, PCDController can be compatible with fine-tuned VDM backbones for versatile downstream tasks. This surprising factor supports domain-specific training, i.e., camera and human motion modules can be trained independently without jointly annotated data. For the global 3D world guidance, we align scenic point clouds (for camera control) and SMPL-X characters (Pavlakos et al., 2019) (for human animation) into the same world-coordinate space via the rigid transformation (Umeyama, 1991), while the 2D keypoints (Xu et al., 2023) bridge the relation of two presentations. Note that our alignment enables complicated motion transfer, covering disparate characters, positions, and viewpoints as verified in the last row of Figure 1.
@@
Extensive experiments validate the efficacy of Uni3C. To evaluate the remarkable generalization of PCDController, we collect an out-of-distribution test set across different domains, where each image has four different camera trajectories. For the joint controllability, we build a comprehensive test set of in-the-wild human videos. GVHMR (Shen et al., 2024) is used to extract SMPL-X as the condition, while three complex and random camera trajectories are assigned for each video. VBench++ (Huang et al., 2024) is employed to verify the overall performance of our method. Uni3C significantly outperforms other competitors, both quantitatively and qualitatively.
@@
2. Related Work
2.1. Camera Control for VDMs
@@
Controlling camera trajectories in video generation has garnered significant attention recently. Some works focused on injecting camera parameters into VDMs to achieve camera controllability (Wang et al., 2024b; He et al., 2025a; Bahmani et al., 2025b, a; Zheng et al., 2025a; He et al., 2025b; Liang et al., 2024), typically utilizing the Plücker ray presentation. For instance, VD3D (Bahmani et al., 2025b) designed a tailored framework for Diffusion Transformer (DiT) (Peebles and Xie, 2023), while AC3D (Bahmani et al., 2025a) further emphasized the generalization with fewer trainable parameters. Moreover, DimensionX (Sun et al., 2024) further decoupled the spatial and temporal control with different LoRAs (Hu et al., 2022). Despite the progress made by these methods, they cannot control the camera movements precisely, particularly when the case is beyond the training domains with an unknown metric scale. Thus, other recent works have employed point cloud conditions via training-based (Yu et al., 2024; Ren et al., 2025; Li et al., 2025b; Feng et al., 2025; Popov et al., 2025) and training-free (Hou et al., 2025; You et al., 2025) manners. However, they fail to accommodate the generalized model design and training strategy to handle imperfect point clouds or out-of-distribution data, especially in motional scenarios involving humans or animals.
2.2. Unified Control for VDMs
@@
Recent works have unified multiple conditions to guide video generation (Wang et al., 2024b; Feng et al., 2025; Geng et al., 2025; Wang et al., 2024a; Chen et al., 2025; Gu et al., 2025; Zheng et al., 2025b). MotionCtrl (Wang et al., 2024b) integrated the camera and object motion controls through separate pose and trajectory injections. Subsequently, researchers further explored the presentation of conditions, such as point trajectories (Feng et al., 2025), point tracking (Geng et al., 2025; Gu et al., 2025), and 3D-aware signals (Chen et al., 2025). Moreover, VidCraft3 (Zheng et al., 2025b) considered the lighting control to VDMs for the first time. For the control of both cameras and human motions, some works focus on the behavior transfer based on camera and human motion estimation (Jiang et al., 2024; Kocabas et al., 2024). Humanvid (Wang et al., 2024a) first unified them in video generation. However, it requires both camera and human guidance from the same source video; otherwise, the human pose would conflict with the camera movements, leading to limited flexibility for the decoupled control. Despite the promising performance demonstrated by these pioneering approaches, they often rely heavily on joint training under various conditions and well-labeled datasets. Furthermore, there has been limited discussion on unified control within foundational VDMs that exceed 10B parameters. Our work offers a solution to address these issues: unifying existing models for different downstream tasks without the need for costly fine-tuning or fully annotated datasets. This strategy is particularly well-suited for large VDMs, enabling models to focus more on enhancing performance within their specific domains.
3. Preliminary: Video Diffusion Models
@@
We briefly review VDMs and Wan2.1 (Wang et al., 2025a) in this section as preliminary knowledge. VDMs are mainly based on the latent diffusion model (Rombach et al., 2022), modeling the conditional distribution $p(z_{0}|c_{txt},c_{img})$, where $z_{0}$ indicate clean video latent features encoded by 3D-VAE; $c_{txt},c_{img}$ denote the text condition for Text-to-Video (T2V) and the optional image condition for Image-to-Video (I2V), respectively. The training of VDM involves reversing the diffusion process by the noise estimator $\epsilon_{\theta}$ as:
(1) $\min_{\theta}\mathbb{E}_{z_{0},t,\epsilon,c_{txt},c_{img}}[\|\epsilon_{\theta}(z_{t},t,c_{txt},c_{img})-\epsilon\|^{2}],$
@@
where $\epsilon\sim\mathcal{N}(0,I)$ indicates Gaussian noise; timestep $t\in[0,T_{max}]$; $z_{t}$ is the intermediate noisy latent state of timestep $t$. Recently, most VDMs have employed Flow Matching (FM) (Lipman et al., 2022) as the improved diffusion process with faster convergence and more stable training. Based on the ordinary differential equations (ODEs), FM formulates the linear interpolation between $z_{0}$ and $z_{1}$, i.e., $z_{t}=tz_{1}+(1-t)z_{0}$, where $t\in[0,1]$ is sampled from the logit-normal distribution. The velocity prediction $v_{\theta}$ is written as:
(2) $\min_{\theta}\mathbb{E}_{z_{0},t,\epsilon,c_{txt},c_{img}}[\|v_{\theta}(z_{t},t,c_{txt},c_{img})-v_{t}\|^{2}],$
@@
where the ground truth velocity denotes $v_{t}=\frac{dz_{t}}{dt}=z_{1}-z_{0}$. Additionally, recent foundational VDMs (Brooks et al., 2024; Yang et al., 2025; Kong et al., 2024; Wang et al., 2025a) are built with DiT (Peebles and Xie, 2023) to achieve more capacities for video generation.
@@
Wan2.1 (Wang et al., 2025a) is an open-released VDM with DiT architecture trained with flow matching (Lipman et al., 2022). umT5 (Chung et al., 2023) is utilized as the multi-language text encoder to inject textual features into Wan2.1 through cross-attention. For image-to-video, Wan-I2V further incorporates features from CLIP’s image encoder (Radford et al., 2021) to improve the results. Uni3C is primarily designed for Wan-I2V with 14B parameters, but we empirically find that it is compatible with the Wan-T2V version as verified in Section 5.4, showing convincing generalization.
4. Method
Figure 2. The overview of Uni3C, which adopts multi-modal conditions. The camera, point clouds, and reference image are assigned to the camera control module called PCDController, while the reference image, SMPL-X (Pavlakos et al., 2019), and Hamer (Pavlakos et al., 2024) are assigned to human animation called RealisDance-DiT (Zhou et al., 2025b).
@@ 4. Method
Figure 3. Pipeline of PCDController, which is built as a lightweight DiT trained from scratch. We first obtain point clouds via monocular depth from the first view. Then, the point clouds are warped and rendered into the video $V_{pcd}$. Input conditions for PCDController comprise rendered $V_{pcd}$, Plücker ray $\mathbf{P}$, and the noisy latent $z_{t}$. Only the PCDController and camera encoder are trainable in our framework. For inference of unified control over camera and human motions, we directly replace the frozen Wan backbone with RealisDance-DiT (Zhou et al., 2025b) without joint fine-tuning.
@@
We first present the overview of Uni3C in Figure 2. Given a reference view $I_{img}\in\mathbb{R}^{3\times h\times w}$, camera trajectories $\{c_{cam}^{i}\}_{i=1}^{N}$ of $N$ target views, and textual condition, Uni3C uses PCDController to produce the target video under specified camera trajectories. This process can be formulated as $p(z_{0}|c_{txt},c_{img},c_{cam})$, where $z_{0}$ indicates the clean latent video feature, $c_{txt},c_{img}$ are textual features from umT5 and image latent condition encoded from $I_{img}$, respectively. We show the pipeline of PCDController in Figure 3 as a core component of Uni3C for generalized camera control (Section 4.1) across downstream tasks. In this work, we focus on human animation (Section 4.2). Subsequently, we introduce the global 3D world guidance illustrated in Figure 5 to unify both camera and human characters into a consistent 3D space for inference (Section 4.3).
4.1. PCDController with 3D Geometric Priors
@@
Following AC3D (Bahmani et al., 2025a), PCDController is designed within a simplified DiT module rather than copying modules and weights from the main backbone as ControlNet (Zhang et al., 2023). To preserve the generalization of Wan-I2V, we follow the insight of training as few parameters as possible once the effective camera control has been achieved. Formally, we reduce the hidden size of PCDController from 5120 to 1024, while zero-initialized linear layers are used to project the hidden size back to 5120 before being added to Wan-I2V. Moreover, as investigated in (Bahmani et al., 2025a; Liang et al., 2024), VDMs mainly determine camera information through shallow layers. Thus, we only inject camera-controlling features into the first 20 layers of Wan-I2V to further simplify the model. Additionally, we discard the textual condition for PCDController to alleviate intractable hallucination and remove all cross-attention modules. In this way, the overall number of trainable parameters for PCDController is reduced to 0.95B, a significant reduction compared to Wan-I2V (14B).
@@
In contrast to merely utilizing Plücker ray as the camera embedding (Bahmani et al., 2025a; Liang et al., 2024), we incorporate much stronger 3D geometric priors to compensate the model simplification, i.e., videos $\{V_{pcd}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 3\times h\times w}$ rendered from point clouds under given camera trajectories. Specifically, we first use Depth-Pro (Bochkovskii et al., 2024) to extract the monocular depth map from the reference view. We then align this depth map into a metric representation using SfM annotations (Schönberger et al., 2016) or multi-view stereo (Cao et al., 2024). Following (Cao et al., 2025), we employ RANSAC to derive the rescale and shift coefficients, preventing the collapse of constant depth outcomes. Subsequently, the point clouds $X_{pcd}\in\mathbb{R}^{hw\times 3}$ are got by unprojecting all 2D pixels from $I_{img}$ into the world coordinate via its metric depth $\hat{D}_{img}$ as follows:
(3) $X_{pcd}(x)\simeq R_{c\rightarrow w}\hat{D}_{img}(x)K^{-1}\hat{x}$,
where $x$ denotes the 2D coordinates of $I_{img}$, while $\hat{x}$ refers to the homogeneous coordinates; $K,R_{c\rightarrow w}$ mean the intrinsic and extrinsic cameras of the reference view, respectively. After that, we render $\{V_{pcd}^{i}\}_{i=2}^{N}$ for the remaining (N1N-1) views by PyTorch3D according to their respective camera intrinsics and extrinsics. Note that the first rendering corresponds to the reference image, i.e., $V_{pcd}^{1}=I_{img}$, to confirm the identity. We apply $V_{pcd}$ to the 3D-VAE of Wan2.1 to achieve $c_{pcd}$ as the point latent condition. To further handle the significant viewpoint changes, which may extend beyond the point clouds’ visibility of the first frame, PCDController also includes Plücker ray embedding (Xu et al., 2024a), $\{\mathbf{P}^{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times 6\times h\times w}$, as the auxiliary condition. $\{\mathbf{P}^{i}\}_{i=1}^{N}$ is encoded by a small camera encoder, comprising causal convolutions and a 4-8-8 downsampling factor to keep the same sequential length as 3D-VAE outputs. The distribution modeling of PCDController is $p(z_{0}|c_{txt},c_{img},c_{pcd},\mathbf{P})$.
Figure 4. Results of PCDController with imperfect point clouds. Benefiting from the well-preserved capacity of VDM, PCDController enjoys robust generation with inferior point clouds.